# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v3

@@ Abstract
Figure 1: Overview of the SCAM datasets and benchmarking framework. A) We create three counterfactual versions of images: SCAM, NoSCAM, and SynthSCAM. B) Evaluation of VLMs via cosine similarity between image and text. C) Prompt-based evaluation of LVLMs.
@@ 1 Introduction
VLMs such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. With the growing use of VLMs in real-world applications, such as medical applications [10, 11, 12] and autonomous systems [13], as well as in Large Vision-Language Models (LVLMs) [14, 15, 16], it is essential to ensure the safety and reliability of these models. However, recent research has revealed that these models are vulnerable to typographic attacks [17, 18, 19], which exploit the reliance on the textual component of multimodal learning.
@@ 3 SCAM Dataset
All images were manually annotated with both object and corresponding attack labels as part of a voluntary community effort. The annotation process was conducted using Label Studio [39]. To standardize the dataset, all images were zero-padded and downscaled to 512×512 pixels using cv2.INTER_AREA [40], an interpolation method that preserves details while minimizing distortion.
@@ 4 Evaluation
We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup. Both evaluations are conducted on the SCAM, NoSCAM and SynthSCAM datasets. The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
@@
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [41, 42, 1, 43, 44]. We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 28, 29, 28] and Llama3.2-vision [45] models. The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API. An overview of selected models is provided in table 1, while LABEL:tab:full-overview presents the complete list of all 106 evaluated models.
@@
Respond strictly with only (a) or (b), nothing else."
@@ 5 Results
Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to as v1). More details about the selection process can be in section A.5.
@@ Appendix A Appendix
A.1 SCAM Examples
@@ A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM, PAINT [38], and RTA-100 [36] datasets.
@@ A.5 Architecture Comparison
Table 10: Models chosen from the OpenCLIP suite for architecture comparison in table 5. The table groups models by architecture category and indicates the corresponding fixed training data. v1 refers to LAION-COCO NLLB [50].

### v4

@@ Abstract
Figure 1: SCAM enables benchmarking safety-critical robustness to typographic attacks in multimodal models. a) We construct three image variants. Real-world attacks in SCAM, a cleaned baseline NoSCAM, and digitally simulated attacks in SynthSCAM allow causal evaluation of typographic vulnerabilities. b) VLMs are evaluated zero-shot by computing cosine similarity between image embeddings and textual labels. Our results show that misleading text embedded in images significantly shifts predictions, indicating overreliance on textual cues. c) LVLMs are assessed using prompt-based classification. Despite advanced architectures, models remain susceptible to typographic attacks in realistic user-facing tasks. Notably, synthetic attacks are as effective as real ones, validating their use for scalable robustness evaluation.
@@ 1 Introduction
Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. With the growing use of VLMs in real-world applications, such as medical applications [10, 11, 12] and autonomous systems [13], as well as in Large Vision-Language Models (LVLMs) [14, 15, 16], it is essential to ensure the safety and reliability of these models. However, recent research has revealed that these models are vulnerable to typographic attacks [17, 18, 19], which exploit the reliance on the textual component of multimodal learning.
@@ 3 SCAM Dataset
[…]
@@ 4 Evaluation
We evaluate VLMs on a zero-shot classification task and LVLMs in a generative prompt-based classification setup. Both evaluations are conducted on the SCAM datasets (SCAM, NoSCAM, SynthSCAM) and additionally on PAINT [38] and RTA-100 [36]. The evaluation aims to demonstrate how typographic attacks, whether real-world or synthetic, affect the accuracy of multimodal foundation models in correctly predicting the object (rather than the attack word) depicted in the image.
@@
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [41, 42, 1, 43, 44]. We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 29, 28] and Llama3.2-vision [45] models. The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API. An overview of selected models is provided in table 1, while LABEL:tab:full-overview presents the complete list of all 106 evaluated models.
@@
Respond strictly with only (a) or (b), nothing else.
@@ 5 Results
Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1). More details about the selection process can be found in section A.5.
@@ Appendix A Appendix
A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
@@ A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP [44] resp. ollama [46] and OpenAI’s API on SCAM, PAINT [38], and RTA-100 [36] datasets. Results on PAINT and RTA-100 are broadly similar to those on SCAM, supporting the consistency of typographic attack effects across datasets.
@@ A.5 Architecture Comparison
Table 10: Models chosen from the OpenCLIP suite for architecture comparison in table 5. The table groups models by architecture category and indicates the corresponding fixed training data. In OpenCLIP, v1 refers to LAION-COCO NLLB [50].